OpenStudy (anonymous):

minimize the function f(x,y,z)=x^2 +y^2 +z^2 subject to the constraints x+2y+3z=6 and x +3y +9z=9

OpenStudy (anonymous):

2x +0y+0z-λ-μ = 0 0x+2y+0z-2λ-3μ = 0 0x+0y+2z-3λ-9μ = 0 x+2y+3z+0λ+0μ = 6 x+3y+9z+0λ+0μ = 9

OpenStudy (anonymous):

so which is the answer? and thank you

OpenStudy (anonymous):

f(x,y,z) = x2+y2+z2